40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Dec 27 05:38:56 2021, rana, Update, Control System, cavity locking loop shape 
    Reply  Mon Jan 24 10:24:49 2022, shruti, Update, Control System, cavity locking loop shape OLTF_fit.pdf
       Reply  Wed Jan 26 17:28:53 2022, shruti, Update, Control System, cavity locking loop shape ctrl_pdh_5kHz_Screenshot.pngctrl_pdh_10kHz_Screenshot.pngscope_red.pdffilter_red.pngctrl_pdh_100kHz_Screenshot.png
          Reply  Mon Jan 31 19:03:43 2022, shruti, Update, Control System, cavity locking loop shape - pomona box TF FilterTF.pdfdata_and_analysis.zip
Message ID: 2841     Entry time: Mon Jan 24 10:24:49 2022     In reply to: 2838     Reply to this: 2843
Author: shruti 
Type: Update 
Category: Control System 
Subject: cavity locking loop shape 

While I played around with locking the cavity I was not able to see the cantilever mode dominate the PDH signal at most settings. The only times I did see it was sometimes when the cavity was losing lock after being locked for a while.

Quote:

Aaron and I previously saw that the 40 Hz cantilever mode was using up most of the PDH signal's linear range while in lock.

This means that the loop has just enough gain to stay locked. Increasing the gain seems to make it unstable, so we want to lower the UGF.

There were some current and temperature settings at which I was not able to lock the cavity with the same overall gain (4.40 on the knob = -6 dB which I did not change) but at the following two settings I was able to lock the cavity with the residual noise being ~30% the PDH pk-pk:

  • I ~ 114 mA, T=8.193 kOhms. PDH pk-pk was 370 mV and the locked residual pk-pk was 130 mV
  • I ~ 130 mA, T=8.356 kOhms. PDH pk-pk was 400 mV and the locked residual was 154 mV

I measured the current using the 'ANALOG OUT' port of the ITC 502 on an oscilloscope using the conversion factor of 20 mA/V.

At the second setting mentioned above I also quickly measured the open loop transfer function similar to elog 2761  and elog 2759 (shown in Attachment 1 and data saved here). Since the gain was lower this time, the UGF was a smaller value of 22 kHz with a larger phase margin.

Attachment 1: OLTF_fit.pdf  28 kB  Uploaded Thu Jan 27 14:07:46 2022  | Hide | Hide all
OLTF_fit.pdf
ELOG V3.1.3-